# Crystal-field magnetostriction of the spin ice under ultrahigh magnetic fields

### v1

@@ Appendix C Elastic constant
$k=25\times\exp[-0.1{(r/a)}^{2}],$ (C1)
where $r$ represents the distance and $a$ denotes the nearest-neighbor bond length. Based on these spring constants, the elastic constant tensor $c^{\alpha\beta}$ can be derived. The resulting tensor closely matches those obtained from sound velocity experiments YNakanishi_2011 and successfully reproduces both the slope of low-energy acoustic phonons as a function of wave vectors and the highest-energy optical phonon mode, which reaches 105 meV in DFT calculations MRuminyDFT_2016 . Equation (C1) produces a table of longitudinal spring constants, which are then converted to an elastic tensor with three independent elements in the units of eV/primitive unit cell: $c_{11}=585.656$, $c_{12}=276.629$, and $c_{44}=276.629$, reflecting the highly symmetric cubic crystal structure. These values are in good agreement with those obtained from ultrasound measurements YNakanishi_2011 , after performing a unit conversion of 1 meV///primitive unit cell volume = 0.000621 GPa, where the primitive unit cell volume is 257.835 Å$\AA^{3}$ Baroudi_2015 .

### v3

@@ Appendix C Elastic constants
$k_{ij}=k_{0}\times\exp(-gr^{2}),k_{0}=800~{}{\rm N/m},g=0.1\AA^{-2}$ (C1)
where $i,j$ denotes spatial directions $x,y,z$, $k_{0}$ and $g$ are adjustable fitting coefficients used to obain the appropriate ealstic constants, and $r$ is bond length in the unit of $\AA$. Springs for neighbours up to a distance of 7.5 $\AA$ are taken into account. Based on these spring constants, the elastic constant tensor $c^{\alpha\beta}$ can be derived. The resulting tensor closely matches those obtained from sound velocity experiments YNakanishi_2011 and successfully reproduces both the slope of low-energy acoustic phonons as a function of wave vectors and the highest-energy optical phonon mode, which reaches 105 meV in DFT calculations MRuminyDFT_2016 . Equation (C1) produces a table of longitudinal spring constants, which are then converted to an elastic tensor with three independent elements in the units of eV/primitive unit cell: $c_{11}=585.656$, $c_{12}=276.629$, and $c_{44}=276.629$, reflecting the highly symmetric cubic crystal structure. These values are in good agreement with those obtained from ultrasound measurements YNakanishi_2011 , after performing a unit conversion of 1 meV///primitive unit cell volume = 0.000621 GPa, where the primitive unit cell volume is 257.835 Å$\AA^{3}$ Baroudi_2015 .